OpenStudy (taepoohwae):

Does 4x-y=0 3x-y=1 have no solution, infinitely many solutions or does it have an ordered pair? If so what's the ordered pair?

OpenStudy (anonymous):

You can use substitution, elimination, or graphing. Substitution: 3x-y=1 -y=-3x y=3x Then you can substitute 3x for y in the other equation. 4x-(3x)=0 4x-3x=0 x=0 Now, substitute 0 for x. 3(0)-y=1 0-y=1 -y=1 y=-1 Thus, the ordered pair is (0,-1)

OpenStudy (taepoohwae):

Thank you

OpenStudy (anonymous):

Welcome :)